In given diatomic gas sound travels at 300m/sec. The root mean square velocity of the molecules of the gas is
a)4m/sec.
b)40m/sec
c)438m/sec
d)43900m/sec

Grade:Select Grade

1 Answers

Aziz Alam IIT Roorkee
8 years ago
[v=\sqrt{\gamma P/\rho }] PM=\rho RT P/\rho =RT/M \Rightarrow v=\sqrt{\gamma RT/M} \Rightarrow RT/M=v^{2}/\gamma V_{rms}=\sqrt{3RT/M} \Rightarrow V_{rms}=\sqrt{3v^{2}/\gamma } Vrms=439.155 m/s